# Identifying compact symmetric objects with high-precision VLBI and Gaia astrometry

### v1

@@ 4.2.3 Variability
Despite limited multi–epoch data, CSOs in our sample exhibit remarkably low variability, fundamentally distinguishing them from more variable radio sources like flat-spectrum radio quasars (FSRQs). While FSRQs commonly show variations exceeding factors of 2–20 due to relativistic beaming effects (e.g. III Zw 2, Wang et al., 2023a), CSOs remain notably steady (Fassnacht & Taylor, 2001; Sotnikova et al., 2024). This stability reflects the dominance of lobe emission that was energized over large time scales, representing the adiabatic expansion and steady growth of radio galaxies (O’Dea, 1998; Orienti, 2016).
@@ 4.2.4 Hotspot advance speeds and kinematical ages
Our kinematic analysis spans over two decades and reveals hotspot separation speeds ranging from to , with most sources exhibiting sub-relativistic speeds. This velocity distribution strongly supports models of young radio source evolution (Kawakatu & Kino, 2006; Bicknell et al., 2018).
We exclude J1111+1955 from our kinematic analysis due to conflicting measurements. Our 2015–2018 VLBI data shows an unusually high apparent speed (), strongly contrasting with previous sub-relativistic measurements () from 1997–2002 (Gugliucci et al., 2005). This discrepancy requires caution, especially as this source resides in a Seyfert II galaxy at (Urry & Padovani, 1995), where large viewing angles typically produce lower apparent speeds. Our limited observation epochs prevent definitive conclusions about this anomaly. Long-term monitoring with consistent temporal sampling is needed to reliably determine its kinematics.
Two high-redshift () quasars (J0943+1702 and J1602+2418) display apparent hotspot separation speeds exceeding , though with considerable uncertainties in their measurements. These CSOs likely represent a subpopulation with mildly relativistic jets oriented at small angles to our line of sight, creating an intermediate class between traditional CSOs and more strongly beamed sources. Their detection at high redshifts may reflect an observational bias, as only the most powerful and possibly more beamed sources are likely to be detected in flux-density-limited samples.
Two sources, J1110+4817 and J1815+6127, show apparent contraction (negative speeds) with bent jet heads, suggesting projection effects where three-dimensional jet paths curve relative to our line of sight (An et al., 2012).
In CSOs with detected radio cores, we observe systematic asymmetries in hotspot advance speeds relative to the core. These asymmetries likely reflect environmental differences on opposite sides of the nucleus (Jeyakumar et al., 2005; Orienti & Dallacasa, 2014), providing direct evidence for the role of host galaxy properties in shaping CSO evolution.
All these exceptional cases reflect the complexity of CSO kinematics and the necessity for long-term monitoring to accurately determine their evolutionary status (Scheuer, 1995; Owsianik & Conway, 1998).
The kinematic measurements reveal CSO ages spanning from 80 to over 1800 years (Table 2). Notably, the youngest sources in our sample are associated with those two high–redshift quasars discussed above, possibly reflecting an observational bias in flux-density-limited samples. The broad distribution of ages suggests that we are observing CSOs at various stages of early evolution (Polatidis & Conway, 2003; Gugliucci et al., 2005; An & Baan, 2012) and supports a scenario where the CSO population comprises two distinct classes: genuine young and growing radio galaxies that will eventually evolve into large-scale radio sources, and older systems where radio lobes are either dying out or frustrated by dense surrounding media.
@@ 4.3.1 Evolutionary Tracks and Classification
The kinematic properties of our well-characterized sample provide strong evidence for distinct evolutionary paths in the CSO population. Our statistical analysis combines 24 CSOs from An & Baan (2012) with 10 additional sources from the current study, all having measured redshifts and hotspot separation speeds (Table 3). The relationships shown in Fig. 2 are derived for this subset of sources, offering insights into the connections between radio power, size, velocity, and age in young radio galaxies.
The power–size correlation (Fig. 2-a) demonstrates a clear dichotomy in the CSO population, supporting our previous proposition (An & Baan, 2012) of two distinct evolutionary tracks. While the predicted radio power would vary with projected source size as , our analysis reveals a clear division at a threshold radio power of approximately W Hz-1 (e.g., Kunert-Bajraszewska et al., 2010). Above this threshold power, CSOs exhibit a relatively tight correlation, suggesting efficient conversion of jet power into source expansion. Below this threshold, sources show a more scattered distribution with limited size growth, indicating their evolution is significantly influenced by their host galaxy environment.
The dichotomy mirrors the established FR I/II division in extended radio galaxies (Fanaroff & Riley, 1974; Hardcastle & Croston, 2020) and reflects theoretical models where high–power jets penetrate the ambient medium while lower–power jets are more susceptible to environmental effects that can impede or redirect their expansion. Although our sample may favor powerful sources due to proper motion measurement requirements, the consistency of this division across multiple observational parameters suggests it represents a fundamental aspect of CSO evolution rather than selection bias.
Figure 2 panels – illustrate systematic differences between high-power (red symbols) and low-power (black symbols) CSOs. High-power sources display systematically larger sizes, higher hotspot advance speeds, and a broader distribution of kinematic ages. Radio powers in panels and follow the theoretical predictions and (where is projected size and is age), indicating increasing efficiency in converting jet kinetic energy to radio emission during the CSO phase as adiabatic losses in the lobes dominate energy dissipation. Additionally, An & Baan (2012) predict an inverse relationship between hotspot velocity and radiative power () for sources with constant power (dashed arrows).
High–power CSOs ( W Hz-1) typically achieve higher hotspot advance speeds , with velocities that tend to increase with source size (Fig. 2-b and 2-c). This behavior characterizes young radio galaxies possessing sufficient energy flux to effectively penetrate their host galaxy environment (Kawakatu & Kino, 2006). In contrast, low-power CSOs show consistently slower speeds () regardless of size, suggesting their expansion is significantly influenced by interaction with the ISM (O’Dea & Saikia, 2021).
The variation of spectral index with both kinematic age and projected source size (Fig. 2 panels and ) reveals a systematic pattern. The youngest and smallest CSOs typically exhibit relatively flat spectra (), indicating strong synchrotron self-absorption in compact regions. As sources age and expand, their spectra systematically steepen, with spectral indices approaching for the oldest and largest CSOs, reflecting the decreasing importance of synchrotron self-absorption as source components expand and energy density decreases.
@@ 4.3.2 Evolutionary Pathways and Feedback Mechanisms
The CSO population exhibits distinct evolutionary pathways corresponding to their power classification. High-power CSOs can maintain sustained growth with jets that effectively penetrate dense ambient media, allowing them to evolve through the complete CSO–MSO–LSO sequence (here MSO and LSO stand for medium-sized and large-sized symmetric objects, respectively). Their broader age distribution indicates longer survival while maintaining double–lobe characteristics. These sources provide mechanical feedback by clearing out nuclear environments and potentially influencing galactic evolution on larger scales (Tadhunter, 2016; O’Dea & Saikia, 2021).
In contrast, low–power CSOs typically have smaller sizes and slower expansion speeds, likely constrained by environments (Baan & An, 2025), potentially preventing expansion beyond the ISM. Their narrower size distribution coupled with broader age distribution suggests periods of stalled or intermittent growth, extending their time in the CSO phase. These sources contribute primarily to localized thermal feedback and ISM regulation. Recent studies of low–power radio galaxies (Baldi, 2023; Chilufya et al., 2024) further suggest that AGN feedback operates through multiple channels depending on jet power, with important implications for galaxy evolution models.
The substantial population of identified lower–power CSOs supports a frustrated or confined evolutionary scenario, where weaker jets interact repeatedly with the dense and clumpy ISM. These sources, while contributing significantly to energy injection in the nuclear regions, may never develop the extended structures characteristic of classical radio galaxies. This distinction between successful and frustrated jets is consistent with recent observations showing many faint radio sources remaining embedded within galactic tori or circum-nuclear star-forming regions. Such sources may primarily influence their immediate nuclear environment rather than affect the broader galactic halo or intergalactic medium, representing a more localized mode of AGN feedback (Hardcastle & Croston, 2020).
Low-power CSOs present a complex evolutionary picture. Their weaker jets often interact with the interstellar medium, leading to external material entrainment and mechanical energy losses (Norman et al., 1982; Kawakatu et al., 2008; Wang et al., 2009). Initially showing CSO morphology, they may rapidly lose defined lobes as jets fragment. These sources potentially connect to radio–weak AGN (Sadler et al., 2002; Croston et al., 2019; Drake et al., 2024), supporting the “frustrated” jet model. They are primarily observable in nearby galaxies such as J0048+3157 (NGC 262) and J2355+4950, and recent observations of radio-quiet quasars reveal weak radio sources with low-power CSO-like morphologies (e.g. Alhosani et al., 2022; Chen et al., 2023a; Wang et al., 2023b).
These failed sources continue to evolve slowly, exerting significant feedback on their host galaxies (Bicknell et al., 2003; Shabala & Alexander, 2009; Mukherjee et al., 2016; Morganti & Oosterloo, 2018; Santoro et al., 2020; Tadhunter et al., 2021; Kukreti et al., 2023). Some high–power CSOs may deviate from their evolutionary path due to changes in AGN activity or environmental interactions, becoming dying CSOs or MSOs (Kunert-Bajraszewska et al., 2010; An & Baan, 2012). These failed CSOs may share observational features with CSO class 2.2, as described by Readhead et al. (2024).
@@ 4.4 Classification and Origin Mechanisms
Recently, Kiehlmann et al. (2024a) proposed a CSO classification scheme similar to our power-based dichotomy, describing an edge-dimmed, low-luminosity class (CSO 1), and an edge-brightened, high-luminosity class (CSO 2). Regarding low-power CSOs (or weak-jet CSOs), we align with the CSO 1 classification of Readhead et al. (2024). These likely constitute the majority of radio galaxies, including numerous young sources only recently identified through low-frequency surveys (e.g., Hardcastle & Croston, 2020; O’Dea & Saikia, 2021).
Readhead et al. (2024) further propose that the transient nature of most high–power CSOs could be explained by ignition in Tidal Disruption Events (TDEs) of giant stars (Rees, 1988). Our analysis suggests that intermittent TDEs are unlikely to be the primary mechanism for most high-power CSOs, supported by several critical factors: (1) The host galaxy demographics differ significantly. TDEs predominantly occur in spiral-type galaxies with pseudobulges (Law-Smith et al., 2017; French et al., 2020), whereas CSOs are typically hosted by luminous elliptical or ultraluminous infrared galaxies with classical bulges and higher central black hole masses (Willett et al., 2010; Tadhunter et al., 2011; Kosmaczewski et al., 2020; Liao & Gu, 2020). The preference for post-merger systems among CSO hosts further distinguishes them from typical TDE hosts. (2) The continuous radiation structures in many CSOs directly contradict expectations for transient events. VLBI observations reveal smoothly connected emission from inner jets to terminal hotspots, with inner jets often displaying higher velocities than the hotspots (Taylor et al., 2000; An et al., 2012). This pattern indicates a continuous supply of energy from the central engine rather than the rapidly decaying energy injection expected from TDEs. (3) The timescale discrepancy is substantial. CSO lifetimes span 20–2000 years based on hotspot kinematic measurements (the present paper), drastically exceeding the typical TDE accretion timescales of months to years (Rees, 1988; Gezari, 2021). Even considering longer–lived TDE disks recently proposed, their maximum duration of years remains orders of magnitude shorter than observed CSO lifetimes (Metzger & Stone, 2016). (4) Energy requirements present a fundamental challenge to the TDE scenario. CSOs require sustained jet powers of erg s-1 over centuries to millennia, corresponding to total energy budgets up to 7 M (Readhead et al., 2024). This vastly exceeds the available energy from typical TDEs (), even considering complete disruption of massive stars (Stone & Metzger, 2016). The required disruption of multiple massive stars in rapid succession lacks observational support and seems physically implausible. (5) Birth rate statistics strongly contradict TDE-driven formation. The CSO birth rate, estimated at Gpc-3 yr-1 (Readhead et al., 1996b, 2024), falls at least two orders of magnitude below the observed TDE rate of Mpc-3 yr-1 (van Velzen 2018; Sazonov et al. 2021). Even considering that only 10% of TDEs might produce relativistic jets (Alexander et al., 2020), the discrepancy remains substantial and difficult to reconcile with the TDE scenario.
However, our interpretation of high-power CSO evolution differs from their view. While Kiehlmann et al. (2024b) suggest that most high-power CSOs (their CSO 2 class) do not evolve into larger jetted AGN throughout their life cycles, characterized by sizes of pc and ages of years, our data suggest these CSOs represent early–stage FR II radio galaxies, with some potentially evolving to kpc–Mpc scales. Some CSOs older than 1000 years and larger than 500 pc show continued growth with edge-brightened lobes, despite moderate hotspot advance speeds (Polatidis & Conway, 2003; An & Baan, 2012).
Moreover, many CSOs exhibit other AGN characteristics that align with long-lived activity rather than transient phenomena. These include prominent emission lines in optical spectra (Gelderman & Whittle, 1994), significant infrared excess indicating dust heating by persistent AGN activity (Tadhunter et al., 2011), and X-ray properties consistent with obscured AGN rather than TDE remnants (Guainazzi et al., 2006; Siemiginowska et al., 2016; Sobolewska et al., 2019). The detection of extended radio emission beyond the CSO scale in some sources (e.g., 0108+388, 1511+0518) provides compelling evidence for recurrent AGN activity spanning much longer timescales than individual TDE events (Baum et al., 1990; Orienti & Dallacasa, 2008; Kunert-Bajraszewska et al., 2010).
A more plausible explanation for CSO energetics involves standard AGN accretion processes driven by gas inflow during galaxy mergers or bar-induced instabilities (Hopkins & Quataert, 2010; Storchi-Bergmann & Schnorr-Müller, 2019). This aligns with the frequent observation of CSOs in merging or interacting systems and explains their significant fraction among infrared-luminous galaxies.
While we cannot rule out that some individual CSOs might be powered by exceptional TDEs, such as the disruption of extremely massive stars by intermediate-mass black holes or unusual multiple-star disruption events, the overall CSO population characteristics strongly favor conventional AGN mechanisms. The CSOs detected by Gaia in our sample may be subject to certain selection biases, particularly favoring optically brighter and lower-redshift sources. However, the radio properties and evolutionary characteristics we observe appear consistent across both Gaia-detected and non-detected populations, suggesting our conclusions about CSO physics remain robust.
The upcoming Square Kilometre Array (SKA) and its precursors will be crucial in detecting weaker CSOs and better understanding sources with low and intermediate powers ( W Hz-1), which may represent a critical transition point between evolutionary pathways. Recent discoveries by SKA pathfinders have enhanced our understanding of CSO evolution through comprehensive flux-density-limited samples of FR I and FR II galaxies (e.g., Hardcastle et al., 2019; Mingo et al., 2019; White et al., 2020; Sejake et al., 2023).
Given their complex evolutionary properties, studying the entire CSO population, rather than focusing only on the brightest or peculiar subsets, is essential for fully understanding extragalactic radio galaxy evolution. Such comprehensive study will likely reveal more complex and diverse evolutionary paths than previously recognized, providing critical insights into both radio galaxy evolution and AGN feedback processes.

### v2

@@ 4.2.2 Variability
Flux density variability in CSOs is generally low (Fassnacht & Taylor, 2001; Sotnikova et al., 2024) compared to blazars and other core-dominated AGN, which often vary by factors of 2–20 due to relativistic beaming effects (e.g., III Zw 2, Wang et al., 2023a). This stability reflects the dominance of lobe emission (O’Dea, 1998; Orienti, 2016) and is a key discriminator from highly variable flat-spectrum sources. For the confirmed CSOs with available multi-epoch observations, their total flux densities across the epochs are presented in Fig. S5. In this work, we quantified whole-source variability as using flux densities from epochs with matched () ranges where feasible. S-band data were prioritised to mitigate short () spacing losses. Typical calibration systematics are folded into the uncertainty via quadratic error propagation. Across the confirmed CSOs, is generally (Table 1), which is consistent with lobe-dominated emission. Individual components can fluctuate more strongly (Table S7, Fig. S5) but do not dominate the integrated flux densities. Apparent outliers are often attributable to mismatch in short () spacing or data-quality issues and are excluded from the quantitative estimate.
The observed size changes in some components correlate with flux-density variations in a pattern consistent with adiabatic expansion of hotspots, where increasing size corresponds to decreasing surface brightness (e.g., Wu et al., 2013). To enhance measurement reliability, we implemented a consistent model-fitting strategy across epochs, using fixed component numbers where appropriate and excluding measurements from epochs with inadequate data quality. The overall stability of CSO emission aligns with theoretical expectations, as their radio emission originates predominantly from extended lobes and hotspots rather than from variable cores, and relativistic beaming effects are minimal due to the large viewing angles typical for these sources.
@@ 4.2.3 Hotspot advance speeds and kinematical ages
| Name | Comp. | PAvar | Age | ||
| (mas yr-1) | (c) | ( yr-1) | (yr) | ||
| CSO | |||||
| J0832+1832 | W-E | 0.0060.004 | 0.060.04 | 0.06 | 1200820 |
| J1110+4817 | SW-NE1 | 0.015 | 0.64 | 0.08 | 830 |
| J1111+1955 | W-E | 0.2950.062 | 5.551.16 | 0.31 | |
| J1158+2450 | NE-C | 0.0010.006 | 0.010.08 | 0.07 | 1160 |
| NW-C | 0.0320.007 | 0.420.09 | 0.060.21 | 8020 | |
| J1234+4753 | SE-C | 0.0020.003 | 0.060.08 | 0.06 | 1840 |
| SE-NW | 0.0040.004 | 0.100.08 | 0.05 | 16801400 | |
| J1244+4048 | SW2-C | 0.0250.045 | 1.102.00 | 0.150.25 | 130 |
| J1247+6723 | SE-NW | 0.0460.010 | 0.330.07 | 0.030.15 | 16030 |
| J1358+4737 | W-E | 0.0210.004 | 0.310.06 | 0.070.10 | 25050 |
| J1407+2827 | W-E | 0.0420.003 | 0.220.02 | 0.04 | 16010 |
| J1511+0518 | W-E | 0.0100.002 | 0.060.01 | 0.040.09 | 480100 |
| J1602+2418 | W-E | 0.0330.006 | 2.540.45 | 0.430.12 | 8010 |
| J1815+6127 | SW-NE1 | 0.033 | 1.16 | 0.44 | 160 |
| NE2-SW | 0.035 | 1.21 | 0.35 | 200 | |
| J1945+7055 | SW2-NE3 | 0.0210.006 | 0.140.04 | 0.02 | 1030270 |
| SW1-NE2 | 0.0130.004 | 0.080.03 | 0.06 | 720220 | |
| SW2-NE2 | 0.0580.006 | 0.380.04 | 0.03 | 28030 | |
Based on two decades of VLBI monitoring, we measured hotspot separation speeds of – (sub-relativistic), consistent with young, expanding sources (Kawakatu & Kino, 2006; Bicknell et al., 2018). We excluded J1111+1955 owing to inconsistent apparent speeds: (2015–2018; our measurement) versus (1997–2002) (Gugliucci et al., 2005). Given its Seyfert II host at (Urry & Padovani, 1995), where large viewing angles disfavour superluminal patterns, and the sparse cadence, we defer interpretation pending uniform, multi-epoch monitoring. Two quasars (J0943+1702, J1602+2418) show separation rates with large uncertainties, plausibly reflecting mildly relativistic jets seen at modest viewing angles. Flux-density-limited selection at high-redshift likely favours such moderately beamed systems. J1110+4817 and J1815+6127 exhibit apparent “contraction” (negative speeds) with bent jet heads, which is consistent with projection and/or pattern-speed effects in curved jets (An et al., 2012). For sources with detected cores, near–far side asymmetries in hotspot speeds likely point to unequal ambient densities on opposite sides of the nucleus (Jeyakumar et al., 2005; Orienti & Dallacasa, 2014), providing a direct probe of host-galaxy properties in shaping CSO evolution. These diversities stress the need for long-term, same-frequency VLBI to mitigate component misidentification and core-shift systematics (Scheuer, 1995; Owsianik & Conway, 1998).
Derived kinematic ages span to yr (Table 2); the apparently youngest objects include the two high- quasars, consistent with flux-limited biases discussed above. The wide age range indicates that we are observing multiple early evolutionary stages (Polatidis & Conway, 2003; Gugliucci et al., 2005; An & Baan, 2012). The data favour a continuum from genuinely young, growing systems to sources slowed or stalled by dense environments.
Kinematic ages (see Table 2) are considered indicators of the order of magnitude. Key caveats are (i) implicit homogeneity and constant-speed assumptions; (ii) asymmetric and clumpy ISM, inducing side-to-side and temporal speed changes; (iii) opacity and core-shift effects when motions are referenced to the core; and (iv) uncertainties in component identification. Despite these limitations, the ages correlate qualitatively with source size (see Section 4.3), supporting the youth scenario while highlighting environmental regulation of growth.
@@ 4.3.1 Evolutionary tracks and classification
Our kinematic analysis combines 24 CSOs from An & Baan (2012) with 8 additional sources from this paper, all having measured redshifts and hotspot speeds (Table 4). Together these 32 objects define the working sample. The relationships in Fig. 2 offer insights into connections between radio power, size, velocity, and age in young radio galaxies.
The power-size correlation (Fig. 2-a) demonstrates clear dichotomy in the CSO population, supporting our two-track evolutionary model (An & Baan, 2012). A practical threshold at W Hz-1 (Kunert-Bajraszewska et al., 2010) separates two regimes: above this power, CSOs show a tight correlation with larger sizes, suggesting efficient conversion of jet power into source expansion; below, sources exhibit a more scattered distribution with limited size growth, indicating environmental influence. This dichotomy mirrors the FR I/II division (Fanaroff & Riley, 1974; Hardcastle & Croston, 2020), reflecting models where high-power jets penetrate ambient medium, while low-power jets face environmental impedance. Although proper-motion requirements are biased in favour of brighter systems, the separation persists across parameters, arguing against a selection artifact.
Self-similar models for young sources expanding into roughly uniform media predict under nearly constant jet power (e.g., Kaiser & Best, 2007); we therefore used – scalings as diagnostics rather than strict self-similar laws. Such models rely on simplifying assumptions including geometric self-similarity, straightforward pressure balance, and quasi-adiabatic magnetic evolution, which may not hold in realistic environments (e.g., Readhead et al., 2024). We therefore treated scaling as diagnostic rather than strict laws. An apparently steeper trend in a subset (sometimes described as ) is empirical and likely shaped by population mixing, environmental diversity, and selection effects. In particular, external pressure gradients can regulate jet collimation and advance speeds (e.g., Baan & An, 2025), potentially dominating over internal jet microphysics. Jets may be collimated by accretion-disk winds, even on sub-parsec scales, which challenges the concept of strict self-similar evolution. From this perspective, the observed trends reflect how jets interact with ambient pressure profiles rather than idealised self-similar growth.
Figure 2 panels – show systematic differences between high-power (red symbols) and low-power (black symbols) CSOs. High-power sources display larger sizes, higher hotspot speeds, and broader age distributions. Radio powers follow theoretical predictions and (where is projected size and is age), which is consistent with increasing dynamical work done against the ambient medium as adiabatic losses come to dominate. High-power CSOs ( W Hz-1) achieve faster hotspot speeds () that increase with source size, characterizing young galaxies with sufficient energy flux to penetrate host environments (Kawakatu & Kino, 2006). Low-power CSOs show consistently slower speeds () regardless of size, suggesting ISM-dominated expansion (O’Dea & Saikia, 2021). Spectral evolution (Fig. 2 panels and ) follows the expected patterns: smaller/youngr CSOs exhibit flatter spectra () due to synchrotron self-absorption, while larger/older sources steepen toward as absorption weakens with expansion and radiative aging accumulates.
We examined relationships among key CSO parameters using Pearson correlation analyses (see Sect. S7). Radio power correlates moderately with both projected size and hotspot velocity (), implying that more powerful jets more effectively drive expansion through the ambient medium, producing larger sources and faster hotspots. Hotspot velocity anti-correlates with kinematic age (), pointing to systematic deceleration as CSOs grow. Projected size correlates positively with spectral index (), consistent with synchrotron aging: larger sources host older electrons that have undergone greater radiative and adiabatic losses, leading to steeper spectra.
@@ 4.3.2 Evolutionary pathways and feedback mechanisms
The CSO population appears to follow power-dependent evolutionary paths. High-power CSOs sustain growth as jets penetrate dense media, evolving along the CSO MSO LSO sequence (MSO/LSO: medium/large-sized symmetric objects). Their broader age distribution implies longer survival with double-lobe morphology, and they provide mechanical feedback that evacuates nuclear gas and can influence host evolution on larger scales (Tadhunter, 2016; O’Dea & Saikia, 2021).
By contrast, low-power CSOs are typically smaller and expand more slowly, likely environment-limited (Baan & An, 2025), often stalling within the ISM. Their narrower size but broader age distributions suggest intermittent or frustrated growth, yielding primarily localized thermal/ISM regulation. Recent work on low-power radio galaxies indicates power-dependent feedback channels with implications for galaxy evolution (Baldi, 2023; Chilufya et al., 2024).
The sizable number of lower power CSOs supports a frustrated/confined scenario in which weak jets repeatedly interact with a clumpy ISM, injecting energy in nuclear regions yet failing to develop large-scale radio structures. This successful versus frustrated dichotomy aligns with observations of faint sources embedded in tori or circumnuclear star-forming regions, implying local rather than halo/IGM-scale feedback (Hardcastle & Croston, 2020).
Low-power CSOs also show entrainment and energy losses that erode clear lobes (Norman et al., 1982; Kawakatu et al., 2008; Wang et al., 2009), potentially linking to radio-weak AGN (Sadler et al., 2002; Croston et al., 2019; Drake et al., 2024). They are most readily observed in nearby hosts (e.g., J0048+3157/NGC 262; J2355+4950), and some radio-quiet quasars show weak, CSO-like morphologies (Alhosani et al., 2022; Chen et al., 2023; Wang et al., 2023b). A fraction of high-power CSOs may still deviate from their evolutionary paths, becoming dying CSOs/MSOs due to nuclear activity changes or environment interactions (Kunert-Bajraszewska et al., 2010; An & Baan, 2012), and may overlap phenomenologically with CSO class 2.2 (Readhead et al., 2024).
@@ 4.4 Classification and origin mechanisms
Recent work proposes a power-based CSO dichotomy similar to ours, with edge-dimmed, low-luminosity (CSO 1) and edge-brightened, high-luminosity (CSO 2) classes (Kiehlmann et al., 2024a). We adopted the low-power/weak-jet class in line with Readhead et al. (2024); such systems likely dominate the radio-galaxy population and include many newly recognised young sources from low-frequency surveys (e.g., Hardcastle & Croston, 2020; O’Dea & Saikia, 2021).
Readhead et al. (2024) argued that many high-power CSOs are transient, powered by TDEs of giant stars (Rees, 1988). Our analysis does not support TDEs as the primary engine for most high-power CSOs on several grounds. Concerning (1) hosts, TDEs prefer spiral galaxies (Law-Smith et al., 2017; French et al., 2020), whereas CSOs typically inhabit luminous elliptic or ultraluminous infrared galaxies with classical bulges and higher BH masses, often in post-merger environments (Willett et al., 2010; Tadhunter et al., 2011; Kosmaczewski et al., 2020; Liao & Gu, 2020). With regard to (2) morphology and continuity, VLBI frequently shows smoothly connected inner jets feeding edge-brightened hotspots, with inner-jet speeds exceeding hotspot speeds (Taylor et al., 2000; An et al., 2012); this implies sustained power, not rapidly fading transients. In terms of (3) timescales, CSO lifetimes of 20–2000 yr (this work) far exceed typical TDE accretion phases (months–years) (Rees, 1988; Gezari, 2021); even longer-lived TDE disks (10 yr) (Metzger & Stone, 2016) remain orders of magnitude too short. Concerning (4) energetics, sustained jet powers of – erg s-1 over centuries to millennia (total up to 7 M; Readhead et al. 2024) greatly exceed typical TDE yields (0.1 M; Stone & Metzger 2016). Multiple massive TDEs in quick succession lack support. Regarding (5) rates, the CSO birth rate (310-5 Gpc-3 yr-1; Readhead et al. 1996b, 2024) is 2 orders of magnitude below the TDE rate (10-5 Mpc-3 yr-1), and even if only 10% produce jets (Alexander et al., 2020), the mismatch persists.
Our interpretation of high-power CSO evolution also differs; whereas Kiehlmann et al. (2024b) posit that most CSO 2s remain compact ( pc, yr), our data indicate many are early stage FR II systems, with some potentially growing to kpc–Mpc scales; several sources with yr and pc show continued edge-brightened growth despite moderate hotspot speeds (Polatidis & Conway, 2003; An & Baan, 2012). Additional AGN-like strong optical lines (Gelderman & Whittle, 1994), infrared excess from dust heating (Tadhunter et al., 2011), and obscured-AGN X-ray spectra (Guainazzi et al., 2006; Siemiginowska et al., 2016; Sobolewska et al., 2019) further support long-lived accretion. Extended radio emission beyond CSO scales in some objects (e.g., 0108+388, 1511+0518) argues for recurrent AGN activity (Baum et al., 1990; Orienti & Dallacasa, 2008; Kunert-Bajraszewska et al., 2010).
A more plausible engine is standard AGN accretion, fuelled by merger-driven inflows or bar instabilities (Hopkins & Quataert, 2010; Storchi-Bergmann & Schnorr-Müller, 2019), which is consistent with the prevalence of interacting hosts and the high fraction among IR-luminous galaxies. While rare CSOs might arise from exceptional TDEs, population-level properties favour conventional AGN. Selection biases from Gaia (optically brighter, lower- systems) are acknowledged, but radio evolutionary trends appear consistent across Gaia-detected and non-detected subsamples.
Forthcoming Square Kilometre Array (SKA) surveys will uncover weaker CSOs and clarify the transition around W Hz-1. SKA pathfinders already improve demographic baselines for FR I/II samples (e.g., Hardcastle et al., 2019; Mingo et al., 2019; White et al., 2020a, b; Sejake et al., 2023). Ultimately, capturing the full CSO population, rather than only bright or peculiar subsets, will reveal more diverse evolutionary paths and sharpen constraints on radio-galaxy growth and AGN feedback.